OpenStudy (anonymous):

Simplify (-1+6i)+(-4+2i)

OpenStudy (anonymous):

Please explain how to do this

OpenStudy (anonymous):

i have \[\sqrt{37}+\sqrt{20}\]

OpenStudy (anonymous):

but i'm not sure it that is correct

OpenStudy (dangerousjesse):

Not quite

OpenStudy (anonymous):

\[\sqrt{57}?\]

OpenStudy (anonymous):

i dont think so. it is just simple -5+8i

OpenStudy (anonymous):

\[\sqrt{89} Then?\]

OpenStudy (dangerousjesse):

Simplify the following: \[-1+6 i-4+2 i \]Evaluate \[-1+6 i-4+2 i. \]\[-1+6 i-4+2 i = (6 i+2 i)+(-1-4) = -5+8 i: \]

OpenStudy (dangerousjesse):

Ayuni, you may want to help them get the answer before just giving it to them :P It's not going to be\[\sqrt{x}\]because you're looking for the simplest version.

OpenStudy (anonymous):

the question just ask u to simplify. then, nothing else u should do after that. just left the answer like that.

OpenStudy (dangerousjesse):

Look through the equation I just wrote and solve for the most simple version.

OpenStudy (anonymous):

owhh.. ok.. thanks 4 that dangeroujesse.

OpenStudy (anonymous):

Oh I see so all you have to do really is combine the imaginary numbers and the read ones?

OpenStudy (dangerousjesse):

Yes, nice job.

OpenStudy (anonymous):

OK thanks :)